OpenStudy (anonymous):

Select the correct graph of -4x + 3y = 6.

OpenStudy (phi):

one way to do these problems is find 2 points on the line, and find the graph that goes through those 2 points.

OpenStudy (phi):

oops. Let me fix my arithmetic!! to find two points, pick simple numbers for x (zero is very simple!) -4x + 3y = 6 replace x with zero: -4 * 0 + 3y = 6 -4 times 0 is 0, and 0 plus 3y is just 3y. you get 3y = 6 divide both sides by 3 to get y=6/3 = 2 all of that means that (0,2) is a point on the line

OpenStudy (phi):

now try x=1 in -4x + 3y = 6 -4 * 1 + 3y = 6 -4 + 3y = 6 add +4 to both sides -4 + 4 + 3y = 6+4 3y = 10 divide by 3 on both sides y = 10/3 = 3 1/3 (at bit ugly!) but (1, 3 1/3) is on the line

OpenStudy (phi):

Can you pick out the graph ?

OpenStudy (phi):

notice that as you go from x=0 to x=1 y goes up from 2 to \( 3 \frac{1}{3}\) that means it slopes up to the right